hi can you help me solve this twp step equation:7=y/8 +-9
Answer by mananth(16946)

7=y/8 +-9
7= y/8-9
add 9
7+9 = y/8 -9+9
15 = y/8
..
multiply by 8
y/8 * 8 = 15*8
y =120
...
m.ananth@hotmail.ca